OpenStudy (anonymous):

What is the common difference in this sequence? 25, 31, 37, 43... A. 25 B. 18 C. 9 D. 6

OpenStudy (anonymous):

D

OpenStudy (anonymous):

different d = a(n) - a(n-1) which means d = a2 - a1 = a3 - a2 ... = 31 - 25 = 37 - 31 .. = 6

OpenStudy (anonymous):

thankyou!